sally is mixing a saline solution with a 9% saline solution to obtain 36 ounces of an 8% saline solution. how many ounces of each of the solution must be used
Answer by stanbon(75887)

sally is mixing a saline solution with a 9% saline solution to obtain 36 ounces of an 8% saline solution. how many ounces of each of the solution must be used
-------
You need to mix water with a 9% solution to get an 8% solution.
--------------------
0.09x + 0*(36-x) = 0.08*36
-----
9x = 8*36
x = 32 ounces (amt of 9% solution)
36-x = 4 ounces (amt of water needed)
================
Cheers,
Stan H.
==============